C alculate the volume of solution A (250ml) that you would need to make 50.00mL of NaOH solution with the same concentration as solution F (0.0001058M). Given that one drop of a liquid is 0.05mL, how many drops of solution A would be needed?
1
Expert's answer
2020-03-25T12:47:00-0400

We will apply the formula

"M_1V_1=M_2V_2"

We are not given the concentration of solution A Let's assume that its concentration is M

substituting the values

"MV=0.0001058\times50\\V=\frac{0.00529}{M}ml"


Since each drop contains 0.05 ml

Then the number of drops contained in V ml = "\frac{0.00529}{0.05\times M}=\frac{0.1058}{M}"